What is the momentum of a vehicle if a force of 1000 N is applied on it in 0.50 seconds if it started from rest?


1
Expert's answer
2022-05-07T14:40:45-0400

The Newton's second law says

"\Delta p=F\Delta t"

Hence, the final momentum

"p_f=p_i+F\Delta t\\
=0+1000*0.5=500\:\rm kg\cdot m/s"